We know how search engines crawl websites: they harvest links from pages and iterate through each to develop an ”image” of the world wide web as a whole. During the last decade of search engine growth, websites have attempted to ”help” or even to more directly influence the crawlers and the subsequent website ranking. For example, conventional wisdom holds that search engines ”prefer” sites that are wide rather than deep, and that having a site index will result in more thorough crawling by the Big Three crawlers: Google, Yahoo, and MSN. But how do crawlers actually behave on websites? Does site design really affect this behavior? We created a series of 10 websites to monitor search engine behavior when crawling with very large websites (wide and deep), as well as their behavior on websites where resources ”disappear.” We analyzed the logs of each of these sites for over a full year to see if the conventional wisdom holds true. GIF animations of Apache log data are used to illustrate the crawling patterns. We found that each search engine exhibited different behavior and crawl persistence, and that site design does appear to affect this behavior. We plot the progress of the crawlers through the sites, and their behaviors regarding the various file types. A side benefit of search engine activity on a site is the ”cached page” which is accessible if the original is unavailable. How long will such pages persist in the cache if the web source page disappears? We examine this issue and the role that the cache and a website’s design play in website preservation.